A 5,5g of a gas contained in 1.50 L at 250C exerts a pressure of 741 torr. What is the molar mass of the gas? (1 atm=760 torr)

1
Expert's answer
2021-11-17T23:45:01-0500

PV"=" nRT


P"=\frac{741}{760}atm=0
.975"


V"=1.5litre"


n"=\frac{5.5}{M}" Where "M=" Molar mass


R"=0.0821atm.L\/mol.K"


T"=25+273.15=298.15K"


"0.975×\frac{5.5}{M}×0.0821×298.15"



M"=\frac{5.5×0.0821×298.15}{0.975×1.5}"



"=92.05g\/mol"